OpenStudy (anonymous):

simplify: (-5x)^3/(25y)^3= a*x^m/b*y^n i know a=-1 M and N =3 but i cant get b

jimthompson5910 (jim_thompson5910):

\[\Large \frac{(-5x)^3}{(25y)^3}\] \[\Large -\frac{125x^3}{1625y^3}\] So a = -125, m = 3, b = 1625, and n = 3

OpenStudy (anonymous):

jim's had a slight typo in the denominator when finding 25^3. Other than that, it's correct.